log2x – 2logx + xlog2
1
Expert's answer
2012-12-13T08:56:48-0500
We can use product and quotient rules for logarithms:
log2x – 2logx + xlog2=log(2x)-log(x^2)+log(2^x)=log( 2x^3*2^x)